OpenStudy (anonymous):

What value of h makes the statement true? (17 + h) + (–9) = 17 + (5 + (–9)) A. –21 B. –5 C. 5 D. 21

OpenStudy (anonymous):

@One098 @satellite73 @SubaruWRX

OpenStudy (shinalcantara):

(17 + h) + (–9) = 17 + (5 + (–9)) simplify first the one at the right side (17 + h) + (–9) = 17 + 5 - 9 (17 + h) + (–9) = 13 transpose 9 to the right (17 + h) = 13 + 9 17 + h = 22 transpose 17 to the right h = 22 -17 h = ?